# LLM2CLIP: Powerful Language Model Unlocks Richer Visual Representation

### v2

@@ Abstract
CLIP is one of the most important multimodal foundational models today, aligning visual and textual signals into a shared feature space using a simple contrastive learning loss on large-scale image-text pairs. What powers CLIP’s capabilities? The rich supervision signals provided by natural language — the carrier of human knowledge — shape a powerful cross-modal representation space. As a result, CLIP supports a variety of tasks, including zero-shot classification, detection, segmentation, and cross-modal retrieval, significantly influencing the entire multimodal domain. However, with the rapid advancements in large language models (LLMs) like GPT-4 and LLaMA, the boundaries of language comprehension and generation are continually being pushed. This raises an intriguing question: can the capabilities of LLMs be harnessed to further improve multimodal representation learning? The potential benefits of incorporating LLMs into CLIP are clear. LLMs’ strong textual understanding can fundamentally improve CLIP’s ability to handle image captions, drastically enhancing its ability to process long and complex texts — a well-known limitation of vanilla CLIP. Moreover, LLMs are trained on a vast corpus of text, possessing open-world knowledge. This allows them to expand on caption information during training, increasing the efficiency of the learning process. However, realizing this potential is challenging. Despite LLMs’ powerful internal comprehension, their autoregressive nature hides this capability within the model, leading to output features with poor discriminability. Our experiments show that directly integrating LLMs into CLIP results in catastrophic performance drops. In this paper, we propose LLM2CLIP, a novel approach that embraces the power of LLMs to unlock CLIP’s potential. By fine-tuning the LLM in the caption space with contrastive learning, we extract its textual capabilities into the output embeddings, significantly improving the output layer’s textual discriminability. We then design an efficient training process where the fine-tuned LLM acts as a powerful teacher for CLIP’s visual encoder. Thanks to the LLM’s presence, we can now incorporate longer and more complex captions without being restricted by vanilla CLIP text encoder’s context window and ability limitations. Our experiments demonstrate that this approach brings substantial improvements in cross-modal tasks. Our method directly boosted the performance of the previously SOTA EVA02 model by 16.5% on both long-text and short-text retrieval tasks, transforming a CLIP model trained solely on English data into a state-of-the-art cross-lingual model. Moreover, when integrated into multimodal training with models like Llava 1.5, it consistently outperformed CLIP across nearly all benchmarks, demonstrating comprehensive performance improvements.
Figure 1: LLM2CLIP Overview. After applying caption contrastive fine-tuning to the LLM, the increased textual discriminability enables more effective CLIP training. We leverage the open-world knowledge and general capabilities of the LLM to better process dense captions, addressing the previous limitations of the pretrained CLIP visual encoder and providing richer, higher-dimensional textual supervision. Experimental results demonstrate that LLM2CLIP can make any SOTA CLIP model even more SOTA ever.
@@ 1 Introduction
CLIP (Radford et al., 2021) is one of the most important multimodal foundational models today. It aligns vision and language signals into a shared feature space by employing a simple contrastive learning loss on large-scale image-text pairs. As a retriever, CLIP supports a wide range of tasks, including zero-shot classification (Qian & Hu, 2024), detection (Lin & Gong, 2023), segmentation (Zhou et al., 2023), and image-text retrieval (Lülf et al., 2024; Koukounas et al., 2024). As a feature extractor, it has become dominant in virtually all cross-modal representation tasks, such as image understanding, video understanding, and text-to-image/video generation. For instance, works like LLaVA (Liu et al., 2023) and Qwen-VL (Bai et al., 2023) leverage CLIP as a feature extractor to obtain visual features for text models, while models like Stable Diffusion (Rombach et al., 2021) and DALL·E 2 (Ramesh et al., 2022) use CLIP’s text encoder to extract textual features for visual models.
What makes CLIP so powerful, particularly as a vision encoder? The core of its strength lies in its unprecedented ability to align visual pretraining with natural language — the carrier of human knowledge. Unlike earlier vision encoders such as VGG and ResNet, which relied on the limited ImageNet dataset and simple image categories with just a few words, CLIP is trained on web-scale data using rich descriptive text. This alignment with language is what sets CLIP apart and unlocks its vast potential. However, since CLIP’s introduction, large language models (LLMs) have advanced significantly. Models like GPT-4 (Achiam et al., 2023) and Llama (Dubey et al., 2024) now demonstrate remarkable language capabilities, yet these advancements have not translated to corresponding improvements in visual representation learning. This prompts the question: can the capabilities of LLMs be harnessed to further improve multimodal representation learning?
@@
In this work, we aim to leverage large language models (LLMs) to enable CLIP to learn more powerful, fine-grained, and rich visual representations. Currently, CLIP is often criticized for its bag-of-words-like perception and the limitations of its text encoder, which suffers from a constrained model size, limited context length, and is trained predominantly on image captioning data, lacking exposure to diverse world corpora. A natural approach would be to replace CLIP’s text encoder with an LLM that embeds rich human knowledge. However, this presents significant challenges. In the cross-modal contrastive learning framework employed by CLIP, the text encoder functions as a set of knowledge anchors in the shared latent space, guiding the alignment of the vision encoder with human knowledge of the physical world. The structure, richness, and discriminability of these knowledge anchors are critical to the visual model’s effectiveness. In contrast, LLMs are primarily designed to predict the next word rather than generate explicit representations of the knowledge they contain. Their textual comprehension abilities and open-world knowledge are latent within the model, rather than present in the output embeddings, making them difficult to utilize in the same explicit manner as CLIP’s text encoder. As a result, using LLMs as a text encoder may not produce linearly separable features, which are crucial for effective feature alignment.
Table 1: Comparison of top-1 Caption Retrieval Accuracy (CRA) for various language models in MS COCO 5K testing set.
@@
Llama3-8B-CC 29.2
Llama3.2-1B-CC 28.0
@@
Figure 2: Real examples of top-1 results from the caption-to-caption retrieval experiment. Before fine-tuning, Llama3’s results were often completely unrelated.
To validate our hypothesis, we designed a caption-to-caption retrieval experiment, as shown in Table 1 and Figure 2. Each image in the MS-COCO dataset has five human-annotated captions. We selected the first two captions as positive samples and performed retrieval across the entire validation set. Using the caption retrieval accuracy (CRA), we evaluated the text model’s ability to differentiate between captions, helping us determine which language model is better suited for CLIP. We found that Llama-3 8B achieved only 18.4% top-1 accuracy, while the standard CLIP-ViT-L reached 66.0% top-1 accuracy. As illustrated in Figure 2, the top-1 caption retrieved by original Llama-3 can be entirely unrelated to the query caption, clearly obstructing effective CLIP learning. Therefore, directly using an LLM to guide CLIP’s visual encoder training is highly constrained.
@@
We believe that enhancing the discriminative power of LLM output tokens through fine-tuning is vital for the success of our proposed approach, allowing the latent capabilities of LLMs to surface. Encouragingly, we found that this can be achieved very efficiently. Specifically, we designed a caption contractive (CC) fine-tuning strategy, applying lightweight fine-tuning to the output tokens of Llama-3 8B using LoRA on the CC3M (Sharma et al., 2018) image captioning dataset. The primary goal of this training task was to adjust the output space, improving the model’s ability to distinguish between different captions. We utilized a supervised SimCSE (Gao et al., 2021; BehnamGhader et al., 2024) contrastive learning loss, where the original captions and re-annotated captions generated by ShareCaptioner (Chen et al., 2023) were treated as positive pairs, pulling them closer. In contrast, all other captions formed a negative sample set that the model learned to push away. Remarkably, after this CC fine-tuning, the caption retrieval accuracy, as shown in Table 1, rose from 18.4% to 73%, which is a 7% improvement over the original CLIP-ViT-L text encoder. This successful fine-tuning process enables us to more effectively harness the open-world capabilities of LLMs for CLIP training.
In a nutshell, we present LLM2CLIP, a novel approach for enhancing visual representation learning through the integration of large language models (LLMs) as shown in Figure 1. This method takes a straightforward yet audacious step by replacing the original CLIP text encoder and augmenting the CLIP visual encoder with the vast knowledge embedded in LLMs. We have identified key obstacles associated with this innovative idea and proposed a cost-effective fine-tuning strategy to overcome them. Our experiments demonstrate that leveraging LLMs as teachers for CLIP training yields substantial improvements, with LLM2CLIP significantly outperforming state-of-the-art pre-trained CLIP models. Our method increased the performance of the previously SOTA EVA02 model by 16.5% on both long-text and short-text retrieval tasks, transforming a CLIP model trained solely on English data into a state-of-the-art cross-lingual model. Furthermore, when incorporated into multimodal model training, such as with Llava 1.5, it consistently achieved comprehensive improvements over EVA02 across nearly all benchmarks. Additionally, the efficient training method proposed by LLM2CLIP ensures that the training cost is nearly identical to fine-tuning the original CLIP. We have also demonstrated that using more powerful language models and larger training datasets can further boost LLM2CLIP’s performance, showcasing the immense potential of our approach. These promising outcomes affirm that we have successfully transformed CLIP into a more general-purpose foundational model. The enhanced LLM2CLIP model possesses richer knowledge and exhibits a remarkable capacity for distinguishing fine-grained and complex long-text semantics. This advancement not only broadens the range of supported downstream tasks but also propels progress across the entire vision domain.
@@ 2 Related Works
Contrastive Language-Image Pre-training (CLIP).
CLIP (Radford et al., 2021) is one of the most important multimodal models today. By training on web-scale data, image captions serve as rich, fine-grained supervision signals for learning representations that align closely with image, freeing CLIP from the limitations of manually defined labels. The richness of the captions endows CLIP with many remarkable capabilities; besides zero-shot image-text retrieval, CLIP features can even achieve region-level correspondence, supporting zero-shot detection and segmentation tasks. Furthermore, CLIP is the most widely used encoder in current image and video understanding tasks. In Tong et al. (2024)’s analysis, it stands out among various visual encoders in MLLMs training as the most significant model, demonstrating its strong capabilities. For generative tasks, CLIP is equally foundational. Its text and image encoders are used to provide text control (Rombach et al., 2021; Ramesh et al., 2022) and image control signals (Blattmann et al., 2023; Zhang et al., 2023) for conditioning generative models. Additionally, the CLIP score (Hessel et al., 2021) has become a crucial metric for assessing the relevance of generated images to text. While the original CLIP only utilized image and text modalities, there have been massive extensions of CLIP to incorporate various other modalities (Dai et al., 2024; He et al., 2023; Moon et al., 2022; Zhang et al., 2022). CLIP definitely has a significant impact on the field of vision and multimodal, and its influence is still growing.
@@
Several works have explored the integration of LLMs into CLIP. JinaCLIP (Koukounas et al., 2024) employed Jina-embeddings-v2 (Günther et al., 2023) as the text encoder, which is a BERT variant with 137M parameters, supporting longer texts. Though achieving similar visual performance to EVA-CLIP (Sun et al., 2023), its text encoder is far behind our used Llama3-8B, limiting the potential benefits from LLMs. T5-V (Jiang et al., 2024) took a different approach by aggregating LLM features at the MLLM layer of LLAVA-next-8B (Li et al., 2024a) while freezing the ViT gradient to focus on adjusting the MLLM output. However, freezing visual encoder does not address the problem that ViT inherently lack complex visual feature extraction capabilities, leading to much worse performance than LLM2CLIP. MATE (Jang et al., 2024) designed a learnable adaptor to bridge the gap between CLIP’s text encoder and LLMs. They trained the CLIP visual encoder using LoRA on a small dataset focused on long-text image retrieval tasks. However, they did not recognize the critical issue we propose in this paper: the poor separability of LLM feature space, which is insufficient for direct support of CLIP training. Our work aims to thoroughly explore the impact of LLMs on CLIP as a foundational model, going beyond the influence on long-text retrieval.
@@
It has been widely recognized that the quality of CLIP’s text embedding is coarse and limited to only 77 tokens. Many works have attempted to extend the length of CLIP captions and retrain CLIP accordingly. For instance, DCI (Urbanek et al., 2024) hired annotators to expand captions based on SAM (Kirillov et al., 2023)’s target hints. LaCLIP (Fan et al., 2024) adopted ChatGPT, Bard, and human rewriting. DreamLIP (Zheng et al., 2024) leveraged ShareCaptioner (Chen et al., 2023) and InstructBLIP to augment 30M captions. Recap-DataComp-1B (Li et al., 2024b) used Llama3-trained LLAVA1.5 to extend 1B captions. In order to handle captions longer than 77 tokens, these methods compromised by summarizing long captions into shorter ones (Urbanek et al., 2024), splitting captions into multiple segments (Fan et al., 2024; Zheng et al., 2024), or finetuning model’s positional encoding to support longer token inputs (Zhang et al., 2024). LLM2CLIP, in comparison, leverages Llama3 as the text encoder, which enables comprehensive understanding of long and dense captions. It not only resolves the limition of token length but also allows for better understanding using LLMs’ open-world knowledge. As a result, our model achieves a breakthrough in performance.
@@ 3 Methods
The contributions of our methodology are threefold: First, we designed experiments to analyze the key reason preventing LLMs from directly participating in multimodal representation learning — the weak discriminability of their output features. Second, we introduced the caption contrastive fine-tuning method, which significantly improves feature discriminability. Third, we developed the LLM2CLIP training framework, which has been proven to be an efficient and effective method for leveraging LLMs to deliver substantial performance improvements to pretrained CLIP models.
3.1 Native LLMs are Ineffective Text Encoders for CLIP
Large Language Models (LLMs), such as Llama-3, are trained on massive world corpora through auto-regression, equipping them with open-world knowledge and enabling them to perform various tasks. We initially hoped to leverage the capabilities of LLMs to directly retrain a CLIP model.
Although LLMs exhibit strong text comprehension, they are difficult to use directly as text embedding models. This is because their knowledge is encapsulated within the model, and their output features are heavily skewed towards individual word predictions. As generative models, they are not trained to ensure good linear separability of output features, making them less effective when used to interpret captions for CLIP. As highlighted by  Chen et al. (2022), cross-modal contrastive learning in CLIP requires that each modality possesses strong internal discriminability.
To evaluate the effectiveness of various language models in terms of text discriminability, and to test whether the native output features of LLMs, as hypothesized, struggle to distinguish image captions, we introduce a new metric: the MS COCO Caption Retrieve Accracy (CRA). MS COCO is a widely used multimodal dataset, containing over 330K images, each with five captions. These captions are written by different human annotators and provide diverse descriptions for each image. In our evaluation on the MS COCO validation set, we use only the first two captions of each image and treat the captions of the same image as positive pairs, while all other captions serve as negative samples. We then perform caption-to-caption retrieval and assess Top-1 accuracy using different language models, defining the result as their CRA score. Higher CRA scores indicate better discriminability of the language models on image captions.
As shown in Table 1, using a pure LLM results in a CRA score of only 18.4%, indicating that the majority of captions cannot be well-separated in the output space. In fact, captions with similar distances may be entirely unrelated, as illustrated in Figure 2. However, the text encoder from the original state-of-the-art CLIP model achieves a CRA score of 66%, proving the inadequacy of native LLM output features in caption discriminability. Consequently, it is challenging to apply LLMs directly in CLIP model training. Subsequent experiments in Table 6 also confirm that replacing CLIP’s text encoder and the corresponding ViT with Llama-3 8B for contrastive learning significantly underperforms the original CLIP.
3.2 A Crucial Lesson: LLM Learning for Image Caption Discrimination
@@
To obtain enough varied descriptions for the same image, we use the ShareCaptioner (Zheng et al., 2024; Chen et al., 2023) modified CC-3M (Sharma et al., 2018) dataset, which provides both original captions and augmented dense captions for each image. These can be treated as positive pairs. We followed the training methodology of LLM2Vec (BehnamGhader et al., 2024), first expanding the LLM’s attention mechanism to bidirectional attention, and employing Masked Next Token Prediction (MNTP) for initialization to achieve better results. Specifically:
First, we transform the LLM’s causal attention mechanism into bidirectional attention, as we no longer need it to retain generative capabilities but rather function as an encoder. Since autoregressive training is no longer required, switching to bidirectional attention improves its ability to capture contextual information. Second, we employ MNTP to train the newly added bidirectional attention mechanism, providing a strong initialization. For a given sequence of $N$ tokens, we mask a subset and predict their values, similar to BERT (Devlin et al., 2018). However, unlike BERT, we adapt this process to fit the nature of LLMs by predicting tokens just before the masked token. We train on both image captions and pure text with equal weighting. In addition to CC-3M, we also use the Wikitext-103 (Merity et al., 2016) dataset to preserve the LLM’s text capabilities and prevent divergence from its original strengths. Finally, we perform the actual caption contrastive fine-tuning, using a supervised SimCSE loss to bring captions of the same image closer together and push apart captions of different images. We use two prompt templates: “Given a caption, retrieve a detailed relevant caption” and “Given a detailed caption, retrieve a short relevant caption,” which are prepended to the query (either original or dense caption) to retrieve the corresponding dense or original caption. Similarly, we use a 1.5M-paired general text dataset curated from  Springer et al. (2024) to maintain strong performance in pure language tasks. All training is efficiently conducted using LoRA and is completed in just one epoch, ensuring low computational cost.
A remarkable result followed: after fine-tuning the LLM as a caption encoder for only one epoch on CC-3M, the CRA score for Llama-3 8B jumped from 18.4% to 73.0%, surpassing the previous state-of-the-art CLIP and EVA models’ text encoders trained on the massive Laion-2B (Schuhmann et al., 2022) and COYO-700M (Byeon et al., 2022) image-text datasets. As shown in Table 6, subsequent experiments demonstrated that after CC fine-tuning, the LLM finally unleashed its powerful capabilities, significantly boosting the performance of the previously state-of-the-art CLIP model, in stark contrast to the results without CC fine-tuning. This breakthrough uncovers the potential of LLMs in CLIP training and removes a major obstacle in leveraging LLMs to advance vision foundation models.
3.3 LLM2CLIP: Makes SOTA CLIP more SOTA ever

### v3

@@ Abstract
CLIP is a foundational multimodal model that aligns image and text features into a shared space using contrastive learning on large-scale image-text pairs. Its strength lies in leveraging natural language as a rich supervisory signal. With the rapid progress of large language models (LLMs), we explore their potential to further enhance CLIP’s multimodal representation learning. This work introduces a fine-tuning approach that integrates LLMs with the pretrained CLIP visual encoder, leveraging LLMs’ advanced text understanding and open-world knowledge to improve CLIP’s ability to process long and complex captions. To address the challenge of LLMs’ autoregressive nature, we propose a caption-to-caption contrastive learning framework to enhance the discriminative power of their outputs. Our method achieves substantial performance gains on various downstream tasks, demonstrating the effectiveness of combining LLMs with CLIP for enhanced multimodal learning.
@@ 1 Introduction
CLIP [32] is one of the most influential multimodal foundational models today. It aligns visual and language signals into a shared feature space by employing a simple contrastive learning loss on large-scale image-text pairs. As a retriever, CLIP supports a wide range of tasks, including zero-shot classification [31], detection [22], segmentation [46], and image-text retrieval [27, 17]. As a feature extractor, it has become dominant in virtually all cross-modal representation tasks, such as image understanding, video understanding, and text-to-image/video generation. For instance, works like LLaVA [24] and Qwen-VL [2] leverage CLIP as a feature extractor to obtain visual features for text models, while models like Stable Diffusion [34] and DALL·E 2 [33] use CLIP’s text encoder to extract textual features for visual models.
Figure 1: Zero-shot multimodal retrieval radar chart. LLM2CLIP significantly improves the performance of original models like CLIP and EVA02 through fine-tuning, substantially outperforming other fine-tuning-based methods such as LongCLIP.
Figure 2: By adjusting the output space of the LLM, we enable the LLM to more effectively act as a private tutor for CLIP, allowing the fine-tuning process of LLM2CLIP to efficiently adjust the pretrained CLIP’s cross-modal space.
Figure 3: LLM2CLIP Overview. After applying caption contrastive fine-tuning to the LLM, the increased textual discriminability enables more effective CLIP training. We leverage the open-world knowledge and general capabilities of the LLM to better process dense captions, addressing the previous limitations of the pretrained CLIP visual encoder and providing richer, higher-dimensional textual supervision.
What makes CLIP particularly powerful as a vision encoder? The core of its strength lies in its unprecedented ability to align visual pretraining with natural language—the carrier of human knowledge. Unlike earlier vision encoders such as VGG and ResNet, which relied on the limited ImageNet dataset and simple image categories with just a few words, CLIP is trained on web-scale data using rich descriptive text. This alignment with language is what sets CLIP apart and unlocks its vast potential. However, since CLIP’s introduction, large language models (LLMs) have advanced significantly. Models like GPT-4 [1] and Llama [8] now demonstrate remarkable language capabilities, yet these advancements have not translated into corresponding improvements in visual representation learning. This raises the question: Can the capabilities of LLMs be harnessed to further improve multimodal representation learning?
@@
In this work, we aim to design a method that incorporates the powerful capabilities of LLMs to fine-tune pretrained state-of-the-art (SOTA) CLIP models, further enhancing their performance. We propose replacing the weak text encoder of CLIP with an LLM, followed by cross-modal contrastive learning training on image-text datasets. While LLMs have already been introduced into CLIP in the InternVL [6] framework, this approach requires large-scale retraining, and the actual performance improvement solely attributable to the LLM remains unclear. In our experiments, we found that directly using LLMs to fine-tune CLIP led to failure in all tasks, with a significant drop in performance across various downstream tasks as in Table 6. We discovered that the output features of LLMs for different texts (or even the same text) were very similar, which led to poor differentiation. This low discriminative power interferes with CLIP’s contrastive learning task. To investigate this further, we designed a caption-to-caption retrieval experiment, as shown in Table 1 and Figure 4. In this experiment, each image in the MS-COCO dataset has five human-annotated captions. We selected the first two captions as positive samples and performed retrieval across the entire test set. Using caption retrieval accuracy (CRA), we evaluated the text model’s ability to differentiate between captions, helping us determine which language model is better suited for CLIP. We found that Llama-3 8B achieved only 5.2% top-1 accuracy, while the standard CLIP-ViT-L reached 25.2% top-1 accuracy. As illustrated in Figure 4, the top-1 caption retrieved by the original Llama-3 model can be entirely unrelated to the query caption, clearly obstructing effective CLIP learning. Therefore, directly using an LLM to guide CLIP’s visual encoder training is highly constrained. Recent work on LLM2Vec also validates this idea, leading us to realize the necessity of fine-tuning the output space of LLMs.
@@
Llama3-8B-CC 29.5
Llama3.2-1B-CC 29.4
Table 1: Comparison of top-1 Caption Retrieval Accuracy (CRA) for various language models on the MS COCO 5K test set.
@@
Figure 4: Real examples of top-1 results from the caption-to-caption retrieval experiment in MS COCO 5K test set. Before fine-tuning, Llama3’s results were often completely unrelated.
@@
We designed a caption contractive (CC) fine-tuning strategy, as Figure 2, pplying lightweight fine-tuning to the output tokens of Llama-3 8B using LoRA on the CC3M [36] image captioning dataset. The primary goal of this training task was to adjust the output space, improving the model’s ability to distinguish between different captions. We utilized a supervised SimCSE [11, 3] contrastive learning loss, where the original captions and re-annotated captions generated by ShareCaptioner [5] were treated as positive pairs, pulling them closer. In contrast, all other captions formed a negative sample set that the model learned to push away. Remarkably, after this CC fine-tuning, the caption retrieval accuracy, as shown in Table 1, rose from 5.2% to 29.5%. This successful fine-tuning process enables us to more effectively harness the capabilities of LLMs for CLIP training.
Building on fine-tuning of LLMs, we designed an efficient approach, LLM2CLIP, to enhance the performance of pretrained CLIP. As in  Figure 3, This method uses an LLM to further fine-tune the pretrained CLIP, allowing the LLM to impart knowledge and capabilities to CLIP’s vision encoder that were lacking during its initial pretraining. LLM2CLIP consists of two stages: in the first stage, we apply caption contrastive (CC) fine-tuning on the LLM with a small dataset to adjust the LLM’s output space. In the second stage, we fine-tune the CLIP vision encoder with the CC-fine-tuned LLM, with LLM gradients frozen, to further improve CLIP’s performance. Experimental results demonstrate that training on only 15M open-source data substantially boosts CLIP’s capabilities, with even 3M data yielding impressive results.
The main contributions of this paper are:
  1. 1.
    We identify the core limitation preventing LLMs from fully leveraging their capabilities within the CLIP training framework—the insufficient separability of LLM output features.
  2. 2.
    We introduce a straightforward approach by applying caption-to-caption contrastive fine-tuning on the LLM’s output layer, enhancing the separability of the output space. This approach shows experimentally that involving LLMs in the CLIP training process can provide substantial benefits.
  3. 3.
    We propose LLM2CLIP, an efficient and effective LLM-assisted fine-tuning method for CLIP, enabling low-cost fine-tuning of pretrained CLIP models and achieving significant performance improvements.
@@ 2 Related Works
Several works have explored the integration of LLMs into CLIP. JinaCLIP [17] employed Jina-embeddings-v2 [13] as the text encoder, which is a BERT variant with 137M parameters, supporting longer texts. Though achieving similar visual performance to EVA-CLIP [38], its text encoder is far behind our used Llama3-8B, limiting the potential benefits from LLMs. E5-V [15] took a different approach by aggregating LLM features at the MLLM layer of LLAVA-next-8B [19] while freezing the ViT gradient to focus on adjusting the MLLM output. However, freezing visual encoder does not address the problem that ViT inherently lack complex visual feature extraction capabilities, leading to much worse performance than LLM2CLIP. MATE [14] designed a learnable adaptor to bridge the gap between CLIP’s text encoder and LLMs. They trained the CLIP visual encoder using LoRA on a small dataset focused on long-text image retrieval tasks. However, they did not recognize the critical issue we propose in this paper: the poor separability of LLM feature space, which is insufficient for direct support of CLIP training. Our work aims to thoroughly explore the impact of LLMs on CLIP as a foundational model, going beyond the influence on long-text retrieval.
@@
It has been widely recognized that the quality of CLIP’s text embedding is coarse and limited to only 77 tokens. Many works have attempted to extend the length of CLIP captions and retrain CLIP accordingly. For instance, DCI [40] hired annotators to expand captions based on SAM [16]’s target hints. LaCLIP [9] adopted ChatGPT, Bard, and human rewriting. DreamLIP [45] leveraged ShareCaptioner [5] and InstructBLIP to augment 30M captions. Recap-DataComp-1B [20] used Llama3-trained LLAVA1.5 to extend 1B captions. In order to handle captions longer than 77 tokens, these methods compromised by summarizing long captions into shorter ones [40], splitting captions into multiple segments [9, 45], or finetuning model’s positional encoding to support longer token inputs [44]. LLM2CLIP, in comparison, leverages Llama3 as the text encoder, which enables comprehensive understanding of long and dense captions. It not only resolves the limition of token length but also allows for better understanding using LLMs’ open-world knowledge. As a result, our model achieves a breakthrough in performance.
@@ 3 Methods
LLM2CLIP is a fine-tuning strategy for CLIP, consisting of two stages. The first stage aims to improve the separability of the LLM’s output space, and the second stage involves fine-tuning by replacing CLIP’s text model with the LLM.
3.1 Caption Contrastive Fine-tuning
@@
To obtain enough varied descriptions for the same image, we use the ShareCaptioner [45, 5] modified CC-3M [36] dataset, which provides both original captions and augmented dense captions for each image. These can be treated as positive pairs. We followed the training methodology of LLM2Vec [3], first expanding the LLM’s attention mechanism to bidirectional attention, and employing Masked Next Token Prediction (MNTP) for initialization to achieve better results. Specifically:
  1. 1.
    Bidirectional Attention Transformation: First, we transform the LLM’s causal attention mechanism into bidirectional attention, as we no longer need it to retain generative capabilities but rather function as an encoder. Since autoregressive training is no longer required, switching to bidirectional attention improves its ability to capture contextual information.
  2. 2.
    Masked Next Token Prediction (MNTP): Second, we employ MNTP to train the newly added bidirectional attention mechanism, providing a strong initialization. For a given sequence of $N$ tokens, we mask a subset and predict their values, similar to BERT [7]. However, unlike BERT, we adapt this process to fit the nature of LLMs by predicting tokens just before the masked token. We train on both image captions and pure text with equal weighting. In addition to CC-3M, we also use the Wikitext-103 [28] dataset to preserve the LLM’s text capabilities and prevent divergence from its original strengths.
  3. 3.
    Caption Contrastive Learning: Finally, we perform the actual caption contrastive fine-tuning, using a supervised SimCSE loss to bring captions of the same image closer together and push apart captions of different images. We use two prompt templates: “Given a caption, retrieve a detailed relevant caption” and “Given a detailed caption, retrieve a short relevant caption,” which are prepended to the query (either original or dense caption) to retrieve the corresponding dense or original caption. Similarly, we use a 1.5M-paired general text dataset curated from [37] to maintain strong performance in pure language tasks.
All training is efficiently conducted using LoRA and is completed in just one epoch, ensuring low computational cost. In our experiments, we found that the most critical step was the third step with the SimCSE loss, which significantly enhanced the caption discriminability of the LLM.
3.2 LLM2CLIP: Use LLM to Finetuning